OpenStudy (anonymous):

Solve using the guess-and-test strategy. Michael and Susan have a combined age of 17 years. Michael is 5 years older than Susan. How old is Michael? A. 5 B. 6 C. 11 D. 22

OpenStudy (anonymous):

11

OpenStudy (anonymous):

c. 11

OpenStudy (debbieg):

Well, since the "combined ages" are 17, you know that his age has to be less than that.

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

let Susan = x years then Michaels = (x + 5) years x + (x + 5) = 17 2x + 5 = 17 2x = 17 - 5 x = 12/2 x = 6 so michael = 6 + 5 = 11